OpenStudy (anonymous):

Algebra help!!!!! Will give someone a medal!!!!! -4x+2y=-11 and -2x+2y=-15 A.) (0.67,-6.83) B.) (2,-9.5) C.) (0.67, -6.83) D.) (-2,-9.5

OpenStudy (anonymous):

D

OpenStudy (anonymous):

Can you help me on some more?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

How many solutions does the system of equations have? Y=5x+7 and y=5x+8 A.) one B.) two C.) none D.) infinitely many

OpenStudy (anonymous):

@kiananichole1010

OpenStudy (anonymous):

C

OpenStudy (anonymous):

@Musics_life

OpenStudy (anonymous):

Can you help me on some ore?

OpenStudy (anonymous):

Can you help me on some more?

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

How many solutions does the system of equations have? X-3y and y=1/3x-5 A.) one B.) two C.) none D.) infinite many

OpenStudy (anonymous):

@kiananicole

OpenStudy (jhannybean):

This is not helping but simply getting answers. and against the code of conduct on OpenStudy,

TheSmartOne (thesmartone):

And read the CoC while you are at it :P http://openstudy.com/code-of-conduct And since you are new here... \(\bf\huge\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8